OpenStudy (anonymous):

if a rectangular area is 64 m^2 and a rope is strung from one corner to the midpoint of one of the two more distant sides, what is the minimum distance of the rope? A. 32m B. 16m C. 6 m D. 4 m

OpenStudy (anonymous):

use pythag theorom

OpenStudy (anonymous):

can someone take a look at mine

OpenStudy (anonymous):

prettychicka45 please go away

OpenStudy (anonymous):

Let: L = lenght H = Height Ar = Area of rectangle At = Area of Triangle d = rope distance from rectangular area: L*H = Ar L*H = 64 The rope will form a triangle with dimensions: Height = h/2 Length = l so the area of the triangle is: At = (l*(h/2))/2 and we also know that At = (1/3)Ar . From these two equations you can solve for h and l With h and l values, use the pythag theorom to find d: \[d = \sqrt{L^{2} + (w/2)^{2}} \]

OpenStudy (anonymous):

but how would I find the minimum?

OpenStudy (anonymous):

the minimun is whatever yu found for d